# onlineReferences

### v5

@@ E.2   Strong restrictions under Fuzzy DID designs
holds.
Here, recall that we define $D_{t}(Z)=\mathbf{1}\{V\geq v_{Zt}\}$ to be the treatment status under $T=t$ in group $Z$.
@@
Assumption 28 implies Assumption 32.
@@
Fix $z\in\{0,1\}$. Then we have $v_{z1}\leq v_{z0}$ or $v_{z1}\geq v_{z0}$.
Therefore, conditioning $Z=z$, we have
[…]
because we assume that the unobserved heterogeneity $V$ is independent of time variable $T$ conditional on the instrument $Z$. ∎
Lemma 1 shows that under Fuzzy DID designs, the treatment status changes between period 00 and period $1$ in a monotone (uniform) way in exposed and unexposed groups, respectively. Note that the treatment participation equation itself does not imply Assumption 32: the independence between the unobserved heterogeneity $V$ and time variable $T$ given $Z$ yields this assumption. Indeed, in this article, we do not assume this restriction.
As we have seen in Subsection 2.4, time also plays the role of instrument in DID-IV settings. From this observation, we can view that in an unexposed group ($z=0$), Assumption 32 corresponds to a monotonicity assumption with respect to time, formally defined below.
Assumption 33 (Monotonicity assumption with respect to time in an unexposed group).
[…]
Assumption 33 requires that time affects the treatment adoption process from period 00 to period $1$ in a monotone (uniform) way in an unexposed group $z=0$. In terms of the notation introduced in Subsection 2.4, this assumption implies that the group variable $G^{T}$ can take three values with non-zero probability. Hereafter, we consider the type of Assumption 33 which rules out the existence of the time-defiers $DF^{T}$ in an unexposed group.
Next, we show that dCDH’s assumptions imply that there are no time compliers $CM^{T}$ and time defiers $DF^{T}$ in an unexposed group ($z=0$).
@@
Suppose Assumption 27, Assumption 28 and Assumption 33 hold. Then, there are no time compliers $CM^{T}$ and time defiers $DF^{T}$ in an unexposed group ($z=0$).
@@
Without loss of generality, suppose that Assumption 33 excludes the time defiers $DF^{T}$ in an unexposed group ($z=0$).
@@
The first equivalence follows from $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 28. Equation (29) implies that there are no time compliers $CM^{T}$ in an unexposed group ($z=0$). ∎
Finally, we show that the condition (27) in Assumption 26 (Fuzzy design) and Assumption 28 imply that in an exposed group ($z=1$), there are no units switching from treatment to non-treatment between period 00 and period $1$. In other words, in Assumption 32, we should have:
$Pr(D_{0}(1)\leq D_{1}(1)|Z=1)=1.$
@@
The condition (27) in Assumption 26 and Assumption 28 imply $Pr(D_{0}(1)\leq D_{1}(1)|Z=1)$=1.
@@
where the first equivalence follows from $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 28. Recall that Assumption 28 implies Assumption 32 from Lemma 1. Suppose that $Pr(D_{0}(1)\geq D_{1}(1)|Z=1)=1$ holds. Then, we have $Pr(D_{1}(1)>D_{0}(1)|Z=1)=0$. This contradicts with the condition (30). ∎
We now clarify how dCDH’s identifying assumptions restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively. Tables 3-4 in Subsection 2.5 show that their identifying assumptions exclude many types in both groups and the restrictions are heavily imposed on an unexposed group: five and three types exist in exposed and unexposed groups respectively, where the types painted in gray color are excluded in both groups.
Comparing Tables 3-4 to Tables 1-2, one can see the difference between dCDH’s restrictions and ours [^29] Tables 1-2 are constructed under panel data settings, but the types represented in these tables remain consistent even when considering repeated cross section settings. The only modification is that as in Tables 3-4, one can only observe either $D_{0}(z)$ or $D_{1}(z)$ for unit $i$ in group $z$ in each table.: in both groups, we only exclude the defiers $DF^{Z}$ by Assumption 15 (Monotonicity), whereas dCDH additionally exclude the type $NT^{Z}\land DF^{T}$ in an exposed group by Lemma 3, and additionally exclude the time defiers $DF^{T}$ and the time compliers $CM^{T}$ in an unexposed group by Lemma 2.
The distinction between dCDH’s restrictions and ours appears to be important in many empirical settings. Unlike the monotonicity assumption, which eliminate the compliers $CM^{Z}$ in both groups, dCDH’s additional restrictions might be difficult to defend in some applications, and the practitioners may want to avoid imposing these restrictions in practice. For instance, in some cases, we may have no valid reason for excluding the specific type $NT^{Z}\land DF^{T}$ in an exposed group, who are not affected by instrument but affected by time. Similarly, in other cases, we may have no justification to exclude the time defiers $DF^{T}$ and time compliers $CM^{T}$ in an unexposed group because of the steady growth of the treatment rate.
E.3   Decomposing the SLATET
In this subsection, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters under Fuzzy DID designs.
The following theorem presents the decomposition result for the SLATET, and clarifies the difference between dCDH’s target parameter (SLATET) and ours (LATET).
Theorem 6.
Suppose Assumptions 26-30 hold. Then, we can decompose the SLATET into two different causal parameters:
$SLATET\equiv E[Y(1)-Y(0)|Z=1,T=1,D_{1}(1)>D_{0}(0)]$
$=E[Y(1)-Y(0)|Z=1,T=1,CM^{T}]Pr(CM^{T}|Z=1,SW)$
$+E[Y(1)-Y(0)|Z=1,T=1,CM^{Z}\land NT^{T}]Pr(CM^{Z}\land NT^{T}|Z=1,SW),$ (31)
where the switcher (SW) in an exposed group ($Z=1$) is the units who become treated in time $T=1$.
@@ E.3   Decomposing the SLATET
The proof directly follows from Table 3 in an exposed group. ∎
Theorem 6 shows that the SLATET can be decomposed into a weighted average of two different causal parameters under Fuzzy DID designs: one parameter measures the treatment effects among the time compliers $CM^{T}$ in an exposed group and the other measures the treatment effects among the type $(CM^{Z}\land NT^{T})$ in an exposed group.
This theorem has a concerning implications; the SLATET may not be policy-relevant even when the instrument is based on the policy change of interest to the researcher. Suppose that the instrument represents an important policy lever for researchers, and they are intrinsically interested in the treatment effects among the compliers $CM^{Z}$ in an exposed group. In this setting, our target parameter, the LATET, is a policy-relevant by construction. However, the SLATET may not be policy-relevant because this parameter captures the treatment effects among the sub-population of the compliers and the time compliers, who are affected by time but not affected by the instrument. As the share of the time compliers $CM^{T}$ increases, the SLATET puts more weight on the treatment effects among the time compliers, which contaminates the interpretation of the magnitude of this paramter even in such policy-relevant contexts.
Using the decomposition result for the SLATET, we finally clarify the interpretation of Assumption 31. We show that this assumption cannot be interpreted as requiring the homogeneous treatment effect among the same sub-population between exposed and unexposed groups.
If we do not assume Assumption 27, we have Table 5 in an unexposed group. [^30] Here, we assume that Assumption 28 and 33 exclude the defiers $DF^{Z}$ and the time defiers $DF^{T}$ in an unexposed group, respectively.
@@
Then, from Table 5, we can interpret the SLATET’ (see Assumption 31) as the treatment effects among the time compliers in an unexposed group in period $1$:
$SLATET^{\prime}$ $\equiv E[Y(1)-Y(0)|Z=0,T=1,D_{1}(0)\neq D_{0}(0)]$
$=E[Y(1)-Y(0)|Z=0,T=1,CM^{T}].$ (32)
@@
By combining equation (31) with (32), one can see that Assumption 31 ($STATET=SLATET^{\prime}$) does not require the homogeneous treatment effects in the same sub-population between the two groups. Rather, this assumption requires that a weighted average of the treatment effects among the time compliers $CM^{T}$ and the type $(CM^{Z}\land NT^{T})$ in an exposed group should be equal to the treatment effect among the time compliers $CM^{T}$ in an unexposed group. This interpretation implies that the requirement of the homogeneous treatment effect in Assumption 31 is less clear than what one might expect.
E.4   Revisiting the issue regarding the use of the Wald-DID estimand
Finally, we revisit the issue pointed by dCDH regarding the use of the Wald-DID estimand, and show that this problem is not serious than one might expect, and does not arise under our DID-IV designs. We also clarify why dCDH exclude so many types in exposed and unexposed groups as seen in Tables 3-4. Hereafter, we assume Assumptions 26-30 in Fuzzy DID designs.
dCDH argue that the Wald-DID estimand requires Assumption 30 to identify the SLATET, and propose alternative estimands that do not rely on this assumption. We first show that their argument depends on the parallel trends assumption in DID designs. Next, we argue that this assumption would be difficult to justify in DID-IV settings.
First, we clarify the interpretation of Assumption 30 by introducing the additional notation. This assumption is actually conditional on $D_{0}(Z)=1$ in addition to $Z$ and $T$, but it is not immediately clear from dCDH why this strong restriction is required only for the sub-population. For each time group $g^{t}$, let $w_{g^{t},z}=Pr(G^{T}=g^{t}|Z=z)$ denote the population share conditional on group $Z=z$, and let $w_{g^{t},z,D_{0}(0)=1}=Pr(G^{T}=g^{t}|Z=z,D_{0}(0)=1)$ denote the population share conditional on $D_{0}(0)=1$ in addition to group $Z=z$. Additionally, let $\delta_{z,t,g^{t}}=E[Y(1)-Y(0)|Z=z,T=t,G^{T}=g^{t}]$ denote the average treatment effect conditional on group $Z=z$, time $T=t$, and time group $G^{T}=g^{t}$.
Then, from Tables 3-4, we can decompose Assumption 30 as follows:
$E[Y(1)-Y(0)|Z=0,T=1,D_{0}(0)=1]$
$=$ $E[Y(1)-Y(0)|Z=0,T=0,D_{0}(0)=1]$
$\iff$ $\delta_{0,1,AT^{T}}=\delta_{0,0,AT^{T}},$ (33)
$E[Y(1)-Y(0)|Z=1,T=1,D_{0}(1)=1]$
$=$ $E[Y(1)-Y(0)|Z=1,T=0,D_{0}(1)=1]$
$\iff$ $\delta_{1,1,AT^{T}}\cdot w_{AT^{T},1,D_{0}(0)=1}+\delta_{1,1,DF^{T}}\cdot w_{DF^{T},1,D_{0}(0)=1}$
$=$ $\delta_{1,0,AT^{T}}\cdot w_{AT^{T},1,D_{0}(0)=1}+\delta_{1,0,DF^{T}}\cdot w_{DF^{T},1,D_{0}(0)=1},$ (34)
where we use the Law of Iterated Expectation, $D_{0}(0)=D_{0}(1)$ (implied by the sharp assignment of the instrument), and $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 28. Equation (33) shows that in an unexposed group ($z=0$), Assumption 30 requires the stable treatment effect for the time always-takers $AT^{T}$. Equation (34) shows that in an exposed group ($z=1$), this assumption requires that a weighted average of the treatment effects among the time defiers $DF^{T}$ and the time always takers $AT^{T}$ is stable during the two periods.
Conditions (33) and (34) are required for Fuzzy DID designs because dCDH rely on the parallel trends assumption in the untreated outcome. In DID-IV settings, we can not observe the average time trends of untreated potential outcomes for the time always takers $AT^{T}$ in an unexposed group and the time always takers $AT^{T}$ and the time defiers $DF^{T}$ in an exposed group because they always adopt the treatment during the two periods. Therefore, if we assume the parallel trends assumption in the untreated outcomes, we should recover them by using conditions (33) and (34).
To see this formally, we use our notation to briefly sketch dCDH’s proof about how the Wald-DID estimand identifies the SLATET under fuzzy DID designs. First, consider the conditional expectation $E[Y|Z=0,T=1]-E[Y|Z=0,T=0]$, which appears in the numerator of the Wald-DID estimand. We can decompose this quantity into two terms.

### v6

@@ D.2.1 Preparation
$D_{0}(0)=1,D_{1}(1)=1$ $AT^{Z}\land AT^{T}$ $CM^{Z}\land DF^{T}$
$D_{0}(0)=1,D_{1}(1)=0$ $DF^{Z}\land AT^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(1)=1$ $AT^{Z}\land CM^{T}$ $CM^{Z}\land NT^{T}$
$D_{0}(0)=0,D_{1}(1)=0$ $DF^{Z}\land CM^{T}$ $NT^{Z}\land NT^{T}$
Table 8: Unexposed group ($z=0$)
observed counterfactual
[…]
$D_{0}(0)=1,D_{1}(0)=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
$D_{0}(0)=1,D_{1}(0)=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(0)=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
$D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
Notes: These tables represent mutually exclusive and exhaustive types under the DID-IV setting.
In the following, we examine which types are excluded by the identifying assumptions under DID-IV and Fuzzy DID, respectively.
D.2.2 DID-IV
In DID-IV designs, the only restriction on treatment adoption behavior is the monotonicity assumption (Assumption 20), which rules out either $DF^{Z}$ or $CM^{Z}$. Therefore, if we rule out $DF^{Z}$, we obtain the following table.
Table 9: Exposed group ($z=1$)
observed counterfactual
[…]
$D_{0}(0)=1,D_{1}(1)=1$ $AT^{Z}\land AT^{T}$ $CM^{Z}\land DF^{T}$
$D_{0}(0)=1,D_{1}(1)=0$ $DF^{Z}\land AT^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(1)=1$ $AT^{Z}\land CM^{T}$ $CM^{Z}\land NT^{T}$
$D_{0}(0)=0,D_{1}(1)=0$ $DF^{Z}\land CM^{T}$ $NT^{Z}\land NT^{T}$
Table 10: Unexposed group ($z=0$)
observed counterfactual
[…]
$D_{0}(0)=1,D_{1}(0)=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
$D_{0}(0)=1,D_{1}(0)=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(0)=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
$D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
Notes: These tables represent mutually exclusive and exhaustive types under DID-IV designs. The types painted in gray color are excluded by Assumption 20.
D.2.3 Fuzzy DID
Next, we consider the treatment adoption behavior under Fuzzy DID. We first prove some lemmas implied by dCDH’s identifying assumptions. Based on these lemmas, we then construct tables that summarize all latent types in the exposed and unexposed groups under Fuzzy DID.
We begin with Lemma 1 below.
@@ D.2.3 Fuzzy DID
Assumption 33 implies the following two monotonicity conditions:
Assumption 37 (IV monotonicity at time $T=1$).
$Pr(D_{1}(1)\geq D_{1}(0))=1\hskip 5.69054pt\text{or}\hskip 5.69054ptPr(D_{1}(1)\leq D_{1}(0))=1.$
Assumption 38 (Time monotonicity for each $z\in\{0,1\}$).
$Pr(D_{0}(z)\geq D_{1}(z))=1\hskip 5.69054pt\text{or}\hskip 5.69054ptPr(D_{0}(z)\leq D_{1}(z))=1$
for each $z\in\{0,1\}$.
@@
We show that Assumption 33 implies Assumption 37. Note that we should have $v_{11}\geq v_{01}$ or $v_{11}\leq v_{01}$. Without loss of generality, suppose $v_{11}\geq v_{01}$. Then, we obtain $D_{1}(1)=\mathbf{1}\{V\geq v_{11}\}\leq D_{1}(0)=\mathbf{1}\{V\geq v_{01}\}$ almost surely.
By comparing $v_{z0}$ and $v_{z1}$ for each $z\in\{0,1\}$, we can also show that Assumption 33 implies Assumption 38. ∎
Next, we prove Lemma 2 below.
@@
Suppose Assumptions 32 and 33 hold. Then, there are no time compliers $CM^{T}$ and time defiers $DF^{T}$ in an unexposed group $Z=0$.
@@
Note that Assumption 33 also implies the time monotonicity given $Z=0$:
$Pr(D_{0}(0)\geq D_{1}(0)|Z=0)=1\hskip 5.69054pt\text{or}\hskip 5.69054ptPr(D_{0}(0)\leq D_{1}(0)|Z=0)=1.$
This implies that at most one of the two switching types ($CM^{T}$ or $DF^{T}$) can exist in an unexposed group $Z=0$. Without loss of generality, suppose that $DF^{T}$ is ruled out in the unexposed group.
@@
The first equivalence follows from $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 33. Equation (22) shows that there are no time compliers $CM^{T}$ in an unexposed group $Z=0$. ∎
Next, we prove Lemma 3 below. This lemma shows that in an exposed group $Z=1$, there are no units switching from treatment to non-treatment between period 0 and period $1$. This lemma also implies that the condition (20) in Assumption 31 and Assumption 33 ensures the existence of the switchers in an exposed group (that is, the SLATET is well defined).
@@
The condition (20) in Assumption 31 and Assumption 33 imply $Pr(D_{0}(0)\leq D_{1}(1)|Z=1)$=1 and $Pr(D_{0}(0)<D_{1}(1)|Z=1)>0$.
@@
Note that Assumption 33 also implies the time monotonicity given $Z=1$:
$Pr(D_{0}(1)\geq D_{1}(1)|Z=1)=1\hskip 5.69054pt\text{or}\hskip 5.69054ptPr(D_{0}(1)\leq D_{1}(1)|Z=1)=1.$
In what follows, we replace $D_{0}(1)$ with $D_{0}(0)$.
@@
where the first equivalence follows from $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 33 and $D_{0}(1)=D_{0}(0)$. Suppose that $Pr(D_{0}(0)\geq D_{1}(1)|Z=1)=1$ holds. Then, we have $Pr(D_{1}(1)>D_{0}(0)|Z=1)=0$. This contradicts with the condition (23). Therefore, we should have $Pr(D_{0}(0)\leq D_{1}(1)|Z=1)=1$. From the condition (23), we also have $Pr(D_{0}(0)<D_{1}(1)|Z=1)>0$. ∎
Note that this lemma also implies that we cannot impose Assumptions 37 and 38 that rule out both $CM^{Z}$ and $CM^{T}$, which contradicts with $Pr(D_{0}(0)<D_{1}(1)|Z=1)>0$.
Finally, we prove Lemma 4 below.
Lemma 4.
The condition (20) in Assumption 31 and Assumption 33 imply $Pr(D_{0}(0)\leq D_{1}(1)|Z=0)$=1. Equivalently, the type $(DF^{Z}\land AT^{T})$ is ruled out in an unexposed group.
@@
From Lemma 3, we have $Pr(D_{0}(0)\leq D_{1}(1)|Z=1)$=1. Under the threshold representation $D_{0}(0)=\mathbf{1}\{V\geq v_{00}\}$ and $D_{1}(1)=\mathbf{1}\{V\geq v_{11}\}$, this implies $v_{11}\leq v_{00}$. Hence, $D_{0}(0)\leq D_{1}(1)$ holds almost surely, and therefore $Pr(D_{0}(0)\leq D_{1}(1)\mid Z=0)=1$ as well. ∎
We now clarify how dCDH’s identifying assumptions restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively. Tables 11-12 show the heterogeneity in treatment adoption behavior under Fuzzy DID. Here, $DF^{Z}$ and $DF^{T}$ are excluded by Assumptions 37 and 38 (under $z=0$), respectively. Since these restrictions depend on the direction of the two monotonicity conditions, we paint these types by blue. $DF^{Z}\land AT^{T}$ and $NT^{Z}\land DF^{T}$ in an exposed group are excluded by Lemma 3 and $DF^{T}$, $CM^{T}$ and $DF^{Z}\land AT^{T}$ in an unexposed group are excluded by Lemma 2 and Lemma 4. Since these restrictions do not depend on the direction of the two monotonicity conditions, we paint these types by gray.
These tables show that four types are excluded in an exposed group and five types are excluded in an unexposed group under Fuzzy DID. Note that if we exclude $DF^{Z}$ and $CM^{T}$ by Assumption 37 and Assumption 38 (under $z=0$), the same number of types are excluded in exposed and unexposed groups, respectively. If we exclude $CM^{Z}$ and $DF^{T}$ by Assumptions 37 and 38 (under $z=0$), four types are excluded in an exposed group and six types are excluded in an unexposed group.
Table 11: Exposed group ($z=1$)
observed counterfactual
[…]
$D_{0}(0)=1,D_{1}(1)=1$ $AT^{Z}\land AT^{T}$ $CM^{Z}\land DF^{T}$
$D_{0}(0)=1,D_{1}(1)=0$ $DF^{Z}\land AT^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(1)=1$ $AT^{Z}\land CM^{T}$ $CM^{Z}\land NT^{T}$
$D_{0}(0)=0,D_{1}(1)=0$ $DF^{Z}\land CM^{T}$ $NT^{Z}\land NT^{T}$
@@
Notes: These tables represent mutually exclusive and exhaustive types under Fuzzy DID designs. The types painted in gray color are excluded by Lemmas 2 and 3. The types painted in blue color are excluded by the two monotonicity assumptions 37 and 38 (under $z=0$).
D.2.4 Comparison
By comparing Tables 9-10 with Tables 11-12, we obtain two implications. First, the restrictions imposed under Fuzzy DID are stronger than those under DID-IV. Second, while the restrictions under DID-IV are symmetric between exposed and unexposed groups, they are asymmetric under Fuzzy DID.
Note that these differences are driven by differences in the identifying assumptions between the two designs. Assumption 33, which is stronger than Assumption 20, imposes time monotonicity and thus rules out one of the two time-switching types ($CM^{T}$ or $DF^{T}$). The condition (20) in Assumption 31, which DID-IV does not impose, rules out the type $NT^{Z}\land DF^{T}$ in the exposed group and the type $DF^{Z}\land AT^{T}$ in the unexposed group. Finally, Assumption 32, combined with Assumption 20, eliminates both $DF^{T}$ and $CM^{T}$ in the unexposed group under Fuzzy DID.
D.3  The difference in target parameter
In this subsection, building on Tables 11-12, we next clarify the difference between dCDH’s target parameter (SLATET) and ours (LATET). Specifically, we decompose the SLATET under the identifying assumptions in Fuzzy DID and discuss its implications.
We prove Theorem 6 below.
Theorem 6.
Suppose Assumptions 31-35 hold. Then, we can decompose the SLATET as follows, which depends on the direction of the two monotonicity assumptions 37 and 38 (under $z=0$).
  • (i)
    Under the monotonicity assumptions that exclude $DF^{Z}$ and $DF^{T}$, we have:
    $SLATET\equiv E[Y(1)-Y(0)|Z=1,T=1,D_{0}(0)<D_{1}(1)]$
    $=E[Y(1)-Y(0)|Z=1,T=1,AT^{Z}\land CM^{T}]Pr(AT^{Z}\land CM^{T}|Z=1,SW)$
    $+E[Y(1)-Y(0)|Z=1,T=1,CM^{Z}\land NT^{T}]Pr(CM^{Z}\land NT^{T}|Z=1,SW).$ (24)
  • (ii)
    Under the monotonicity assumptions that exclude $DF^{Z}$ and $CM^{T}$, we have:
    $SLATET$ $\equiv E[Y(1)-Y(0)|Z=1,T=1,D_{0}(0)<D_{1}(1)]$
    $=E[Y(1)-Y(0)|Z=1,T=1,CM^{Z}\land NT^{T}].$
  • (iii)
    Under the monotonicity assumptions that exclude $CM^{Z}$ and $DF^{T}$, we have:
    $SLATET$ $\equiv E[Y(1)-Y(0)|Z=1,T=1,D_{0}(0)<D_{1}(1)]$
    $=E[Y(1)-Y(0)|Z=1,T=1,AT^{Z}\land CM^{T}].$
Proof.
Regarding the case (i), the proof directly follows from Table 11. We can also construct the table in an exposed group for the cases (ii) and (iii), which ensures the above interpretation of the SLATET. ∎
Theorem 6 has several important implications. First, the interpretation of the SLATET depends on the direction of the two monotonicity conditions 37 and 38 (under $z=0$). This implies that, if empirical researchers wish to understand which latent type’s causal effect is identified by the SLATET, they need to specify the direction of these monotonicity conditions ex ante.
Second, the SLATET may fail to be a policy-relevant treatment parameter (Heckman and Vytlacil (2001)), even when the instrument represents the policy change of interest to the researcher. Specifically, under case (i), the SLATET is contaminated by the treatment effects among the type $AT^{Z}\land CM^{T}$, who are affected by time but not affected by the instrument. Under case (iii), the SLATET captures only the treatment effects among the type $AT^{Z}\land CM^{T}$.
Third, the SLATET may be defined on a strictly smaller population than the LATET. In particular, under case (ii), the SLATET identifies the causal effect among the type $CM^{Z}\land NT^{T}$, which is a strict subpopulation of $CM^{Z}$. This implies that even when empirical researchers choose the direction of the two monotonicity conditions so that the SLATET is policy relevant, it will measure the treatment effects for a narrower population than the LATET.
Remark 4.
Using the decomposition result for the SLATET, we can also clarify the interpretation of Assumption 36, which dCDH introduce as an alternative to Assumption 32.
First, we describe the heterogeneity in treatment adoption behavior in the unexposed group without Assumption 32. If we do not impose Assumption 32, we obtain Table 13 for the unexposed group. Here, we consider the case in which Assumptions 37 and 38 exclude $DF^{Z}$ and $DF^{T}$ in the unexposed group, respectively.
Table 13: Unexposed group ($z=0$)
observed counterfactual
[…]
$D_{0}(0)=1,D_{1}(0)=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
$D_{0}(0)=1,D_{1}(0)=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(0)=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
$D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
Notes: This table represents mutually exclusive and exhaustive types in an unexposed group under Fuzzy DID without Assumption 32. The types painted in gray color are excluded by Lemma 4. The types painted in blue color are excluded by the two monotonicity assumptions 37 and 38 (under $z=0$).
@@ D.3  The difference in target parameter
Then, from Table 13, we can interpret the SLATET’ (see Assumption 36) as the treatment effects among the type $AT^{Z}\land CM^{T}$ in an unexposed group:
$SLATET^{\prime}$ $\equiv E[Y(1)-Y(0)|Z=0,T=1,D_{0}(0)\neq D_{1}(0)]$
$=E[Y(1)-Y(0)|Z=0,T=1,AT^{Z}\land CM^{T}].$ (25)
@@
Combining (25) with (24), we see that Assumption 36 ($SLATET=SLATET^{\prime}$) does not require homogeneous treatment effects for the same latent type across exposed and unexposed groups. Rather, it equates a weighted average of treatment effects among types $AT^{Z}\land CM^{T}$ and $CM^{Z}\land NT^{T}$ in the exposed group with the treatment effect for type $AT^{Z}\land CM^{T}$ in the unexposed group.
D.4  The use of the Wald-DID estimand
In this section, we explain why the role of the Wald–DID estimand differs between Fuzzy DID and DID-IV designs. In DID-IV, we adopt the Wald–DID estimand as a natural estimand for identifying the LATET. By contrast, dCDH point out that, under Fuzzy DID, the Wald-DID estimand requires the stable treatment effect assumption (Assumption 35) to identify the SLATET. In the following discussion, we show that this difference arises because Fuzzy DID and DID-IV rely on different types of the parallel trends assumption.
D.4.1 Why the Wald-DID estimand requires Assumption 35 under Fuzzy DID
We begin by explaining why the Wald-DID estimand requires Assumption 35 under Fuzzy DID. We show that this requirement does not stem from the Wald-DID estimand itself, but rather from imposing the parallel trends assumption on untreated outcomes.
To make this point precise, we first clarify Assumption 35 using latent treatment adoption types. This assumption is in fact conditional on $D_{0}(Z)=1$ in addition to $Z$ and $T$, but it is not immediately clear from dCDH why such a strong restriction is imposed only on this subpopulation.
From Tables 11-12, we can reinterpret Assumption 35 as follows [^26] Under the identifying assumptions in Fuzzy DID, the type $DF^{Z}\land AT^{T}$ is eliminated in both groups. Therefore, we write the type $AT^{Z}\land AT^{T}$ as $AT^{T}$.:
$E[Y(1)-Y(0)|Z=0,T=1,D_{0}(0)=1]$ $=E[Y(1)-Y(0)|Z=0,T=0,D_{0}(0)=1]$
$\iff E[Y(1)-Y(0)|Z=0,T=1,AT^{T}]$ $=E[Y(1)-Y(0)|Z=0,T=0,AT^{T}],$ (26)
$E[Y(1)-Y(0)|Z=1,T=1,D_{0}(1)=1]$ $=E[Y(1)-Y(0)|Z=1,T=0,D_{0}(1)=1]$
$\iff E[Y(1)-Y(0)|Z=1,T=1,AT^{T}]$ $=E[Y(1)-Y(0)|Z=1,T=0,AT^{T}].$ (27)
Here, we use $D_{0}(0)=D_{0}(1)$ from the sharp assignment of the instrument $v_{10}=v_{00}$. Equations (26)-(27) show that Assumption 35 actually requires the stable treatment effect only for the time always-takers $AT^{T}$ within each group.
We now show that Assumption 35 is not required by the Wald-DID estimand per se, but arises because dCDH impose the parallel trends assumption on untreated outcomes. The key point is that imposing the parallel trends assumption in untreated outcomes under the DID-IV setting forces us to impute the counterfactual untreated potential outcome for the time always-takers $AT^{T}$, which in turn requires imposing Assumption 35.
To see this formally, we focus on the numerator of the Wald-DID estimand. Under Fuzzy DID, we can decompose the conditional expectations E[Y|Z=0,T=1]E[Y|Z=0,T=0]E[Y|Z=0,T=1]-E[Y|Z=0,T=0] and E[Y|Z=1,T=1]E[Y|Z=1,T=0]E[Y|Z=1,T=1]-E[Y|Z=1,T=0] as follows.